From GlueXWiki
Jump to: navigation, search

March 21, 2013 FDC meeting


  1. Production Construction Tracking (Dave)
    • Package #3 repair
  2. Installation schedule(Lubomir)
  3. Engineering (Bill)
    • Cooling manifold installation and tests
  4. Electronics (Chris, Nick)
  5. Tests at EEL126 FDC E-log (Beni)
  6. Other


Participants: Fernando, Eugene, Bill, Dave, Chris, Nick, Simon, Beni, and Lubomir.


- Dave: the last two wire frames populated by Chris. The only remaining operation is to deaden four wire planes. The extra spare cathode type-2 is tensioned, but the cards are not installed yet by purpose (for easy storage).

- Package #3 repair: last Friday we started testing the top cell (after we opened it once and removed a lint) and it was tripping again. On Monday we opened the package for a second time and looked more carefully: there were some small balls on two sense wires and a section of a field wire that looked suspicious. We tried to clean these places and closed the package, turned out again without success. Troubleshooting the problem is difficult because it appears from time to time once per several hours. The dark currents are very low, there's no constant noise and we can't localize the problem. We decided to open the package tomorrow morning, put the wire frame on the stringing table at 90deg (Bill's idea) and scan the wires with the camera. We will clean if possible, or replace the bad wires and next week will give it another try. If it still doesn't work, we will put a spare wire frame (#28, it had higher dark (~350nA) current but was stable for weeks).


- The last, "frozen" schedule moved the FDC installation just by a couple of days. Now the FDC insertion is in the middle of January 2014, dictated mainly by the electronics. Since the testing outside of the magnet should start much earlier we must be able to test the different packages connecting to the same electronics.

- With Eugene, Bill, and Fernando together there were long discussions about the cabling. There was a concern (Lubomir) that with the current installation scheme, when the FDC is inserted one needs extra cable that is one cart long. In total we will need about 3 x 2.5 carts which might be close to 18m, the length of the signal cables. Bill will check that with the I-deas model. Fernando insisted that we have to keep the cables fixed to the racks at the downstream end and never un-bundle them. In this case the question is where to put the extra bundles that are 6-7m long after the insertion of the FDC: if possible somewhere at the front of the magnet??

- Installation of the cables from the two bottom quadrants will be difficult because of the small opening going to the bottom (below the platform) trays. Bill had an idea to elevate the racks so that the space below them will be accessible from the front and the cables can run there. This makes the cable installation much easier, but requires taller people to service the electronics. Other options: 1) disconnecting the cables from the electronics before FDC insertion and re-connecting them again after the insertion (Lubomir), but has to be done every time we take FDC out of the magnet. 2) Putting all the cables on the top side of the racks (Eugene), but the length might not be enough and it will be more difficult to work there.


- Bill and Casey tested the cooling of package #4 that has the manifolds installed: first with CO2 at 100psi and then with Flurinert, pressured again with CO2. No significant leakage was found with CO2, but with Flurinert it was leaking like a sieve (Bill) however mostly at the temporary barb connections needed to do the test, while the compression connections didn't show obvious problems. It turned out the leakage detector we have is not very sensitive to Fluorinert. So, for the next test we need better fittings and leakage detector.

- Bill ordered three more manifold (Norel material), because three were damaged during the installation. Bill also ordered new G10 pieces that are needed to hold the tubes at the 210deg mounting place.


- Fernando: we have already 16 fADC125 and Cody is testing them. Some delays in this pre-production and the procurement is helping to accelerate it.

Tests at EEL126

- Beni at the beginning of the meeting showed results from the tests of the last production package (#2): entry 108 in the FDC E-log. All the channels without any exceptions are working fine. One of the HV connectors (from the side of the detector) had some issues, but was connected.

- The plan is to start preparing for the tests with the full electronics. For this tests, Beni's plan (not present at these discussions) is not to read all the samples, but to have the charge integration and timing extraction done in the ROCs. Eugene: what is the advantage of this compared to just writing all the samples to the disk and later do this off-line. Obviously, writing and reading the files will take much more time, but Beni may have better explanation. As for the cooling Bill says it must be ready in few weeks. A water chiller will be borrowed from FEL, instead of repairing the one they offered us before.


- NO FDC meeting next week.